OpenStudy (anonymous):

What is the sum of a 6-term geometric sequence if the first term is 11, the last term is -11,264 and the common ratio is -4?

OpenStudy (cwrw238):

S6 = 11 * (1 - (-4)^6) ---------- 1 - (-4)

OpenStudy (cwrw238):

you don't need to use the last term

OpenStudy (jdoe0001):

\(\bf S_n =a_1\left(\cfrac{1-r^n}{1-r}\right)\\ a_1 = \text{first term}\\ \text{r = ratio}\\ n = n^{th}\text{ term}\)

OpenStudy (jdoe0001):

read cwrw238 above